OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1).

OpenStudy (anonymous):

(y -1) = 1-2/ 2+3 ( x-2) => y -1 = -1/5 (x-2) => 5y - 5 = -x + 2 => x + 5y = 7

OpenStudy (anonymous):

the answers to pick from are... y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −(x − 2) y − 2 = −(x + 3)

OpenStudy (anonymous):

either option or question is incorrect.......... please check.... carefully.............

OpenStudy (anonymous):

I copied it right from the worksheet off the internet that i need to do..

OpenStudy (anonymous):

you can check............. Is the point (2,1) satisfy any one given option??????????

OpenStudy (anonymous):

something wrong!!!!!!!!!!!!!!

OpenStudy (anonymous):

i figured it out.. The answers to pick from are y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −1/5(x − 2) y − 2 = −1/5(x + 3)

OpenStudy (anonymous):

y − 2 = −1/5(x + 3) is correct..............

OpenStudy (anonymous):

THANK YOU! can you help me with a few more?

OpenStudy (anonymous):

u welcome......... u can also contact by sudhanshu.cochin@gmail.com or skype : sudhanshu_kmr